Suppose that y = f(x) : (−∞, ∞) → (−∞, ∞) is infinitely differentiable and has a local minimum at 0. Prove that there exists a disc centered on the y axis which lies above the graph of f and touches the graph at the point (0, f(0)).


1
Expert's answer
2021-05-07T09:35:32-0400

Given function is-

"y=f(x):(-\infty,\infty)\to (-\infty,\infty)"


The given function is differentiable and have local minimum at x=0

So

"f'(x)=0, f(0)=0"


Also x=0 is the point of local minimum ao The value of "f''(x)>0," Hence The region must lies in the positive direction.


Also It exist at "x=0," The centered of the disc must lie on y-axis.


Hence There exist a disc centered on the y-axis which lies above the graph of f and touches the graph at the point (0, f(0)).